OpenStudy (anonymous):

One angle is three times its supplement increased by 40(degrees). Find the measures of the two supplementary angles.

OpenStudy (anonymous):

x=3(180-x)+40 x=540-3x+40 x=540-3x 580/4x = 145 Am I doing something wrong?